OpenStudy (anonymous):

11 *4th root of 9 - 3 cube root of 9 - *4th root of 9 -5 cube root of 9= ?

OpenStudy (anonymous):

\[11\sqrt[4]{9}-3\sqrt[3]{9}-\sqrt[4]{9}-5\sqrt[3]{9}\] \[10\sqrt[4]{9}-8\sqrt[3]{9}\] \[10\sqrt{3}-8\sqrt[3]{9}\]